OpenStudy (anonymous):

square root of 216? prefer steps shown. and how do you make the square root symbol on the computer? hahah

OpenStudy (dumbcow):

" sqrt{ } "

OpenStudy (anonymous):

216 is just 6^3, so the answer is 6 radical 6

myininaya (myininaya):

\[\sqrt[b]{x}\] sqrt[b]{x}

OpenStudy (anonymous):

uh thanks but that doesn't really help..

myininaya (myininaya):

it should that is what you type into the equation thing below to write the square root symbol on the computer

OpenStudy (anonymous):

okay thanks, but i need help with the \[\sqrt{216}\] and an explanation ...

myininaya (myininaya):

\[\sqrt{216}=\sqrt{3 \cdot 72}=\sqrt{3 \cdot 8 \cdot 9}=\sqrt{3 \cdot 2 \cdot 4 \cdot 9}\] \[=\sqrt{9}\sqrt{4}\sqrt{6}=3 \cdot 2 \sqrt{6}=6\sqrt{6}\]

OpenStudy (dumbcow):

chess machine already did that 216 = 6*6*6 sqrt{ 6*6} = 6 -> sqrt{216} = 6*sqrt{6}